# Can Performant LLMs Be Ethical? Quantifying the Impact of Web Crawling Opt-Outs

### v1

@@ Appendix C Matched 50-grams in the compliant dataset
Part of an article quoted on unilad.co.uk (unavailable) []delimited-[][...][ … ] The billionaire writes in the The New York Times: Every one of us — citizens, philanthropists, business and government leaders — should be troubled by the enormous gap between how little of our natural world is currently protected and how much should be protected. It is a gap that we must urgently narrow, before our human footprint consumes the earth’s remaining wild places. []delimited-[][...][ … ]
@@
Republished article on alessandrosicurocomunication.com Google and a corporation associated with NASA are forming a laboratory to study artificial intelligence by means of computers that use the unusual properties of quantum physics. Their quantum computer, which performs complex calculations thousands of times faster than existing supercomputers, is expected to be in active use in the third quarter of this year. The Quantum Artificial Intelligence Lab, as the entity is called, will focus on machine learning, which is the way computers take note of patterns of information to improve their outputs. []delimited-[][...][ … ]
C.2 medicalxpress.com
@@
medicalxpress.com and healthed.com.au republished Stanford University Medical Center People suffering from a debilitating and often discounted disease known as chronic fatigue syndrome may soon have something they’ve been seeking for decades: scientific proof of their ailment. Researchers at the Stanford University School of Medicine have created a blood test that can flag the disease, which currently lacks a standard, reliable diagnostic test. “Too often, this disease is categorized as imaginary,” said Ron Davis, PhD, professor of biochemistry and of genetics. When individuals with chronic fatigue syndrome seek help from a doctor, they may undergo a series of tests that check liver, kidney and heart function, as well as blood and immune cell counts, Davis said. []delimited-[][...][ … ]
@@
medicalxpress.com and hamodia.com (unavailable) republished The Assosciated Press Since late last year, people in the central Chinese city of Wuhan have been infected with a viral pneumonia whose cause was unknown. The outbreak raised the specter of another SARS epidemic, which killed hundreds in 2002 and 2003. A preliminary investigation has now identified the respiratory disease as a new type of coronavirus, Chinese state media reported Thursday, citing scientists handling the investigation. As of Sunday, local authorities reported 59 people with the illness. Seven were in critical condition, while the rest were stable. Eight were discharged Wednesday night after they didn’t exhibit any more symptoms for several days. []delimited-[][...][ … ]
C.3 stackexchange.com
@@
Quote found on www-cs-faculty.stanford.edu matches quoted answer []delimited-[][...][ … ] Many readers are no doubt thinking, ”Why does Knuth replace MIX by another machine instead of just sticking to a high-level programming language? Hardly anybody uses assemblers these days.” Such people are entitled to their opinions, and they need not bother reading the machine-language parts of my books. But the reasons for machine language that I gave in the preface to Volume 1, written in the early 1960s, remain valid today: []delimited-[][...][ … ]
@@
Poetry found on en.wikiquote.org matches quoted answer []delimited-[][...][ … ] And the Raven, never flitting,
@@
That lies floating on the floor, Shall be lifted—nevermore. []delimited-[][...][ … ]
Appendix D Additional experimental results
D.1 Top 50 filtered URLs
@@ D.2 More results
Appendix E NYTimes Memorization Experiments
@@
E.1 Benchmark Creation Prompts
@@
E.2 Additional Results
Table 6 and Table 7 show the LCCS and BLEU scores for 128 and 512 prompt lengths, respectively, across all model configurations. For each prompt length, we filter those articles with less than the prompt length since there is no completion for this kind of prefix. You can also see the evaluation results on the 1T token pretrained model from section 5.3 in Table 5. For all the generations, we used temperature 0.0 (greedy decoding).
To verify the validity of the NYTimes-MCQ, we performed a sanity check using DeepSeek-R1. When the model was supplied with the relevant article at inference, it attained a 99.0% accuracy. When the article was withheld, the model achieved 91.7% accuracy. For additional comparison, we tested GPT-4o on NYTimes-MCQ and observed 89.8% accuracy.

### v2

@@ Appendix D Matched 50-grams in the compliant dataset
Part of an article quoted on unilad.co.uk (unavailable) [][...] The billionaire writes in the The New York Times: Every one of us — citizens, philanthropists, business and government leaders — should be troubled by the enormous gap between how little of our natural world is currently protected and how much should be protected. It is a gap that we must urgently narrow, before our human footprint consumes the earth’s remaining wild places. [][...]
@@
Republished article on alessandrosicurocomunication.com Google and a corporation associated with NASA are forming a laboratory to study artificial intelligence by means of computers that use the unusual properties of quantum physics. Their quantum computer, which performs complex calculations thousands of times faster than existing supercomputers, is expected to be in active use in the third quarter of this year. The Quantum Artificial Intelligence Lab, as the entity is called, will focus on machine learning, which is the way computers take note of patterns of information to improve their outputs. [][...]
Republished article on bike.enginerve.com In a post today on the NYTimes there was an interesting piece on the nature of cycling and head injuries. I still object to laws requiring a helmet to be worn at all times and wear my helmet constantly for touring and commuting. Assumption of risk, and cost, is an issue and not one to be lightly avoided. And neither are the responsibility to utilize appropriate safety gear. Cycling Is the Top Sport for Head Injuries Anahad O’Connor tackles health myths in this NY Times Post. Last week, New York City began its long-awaited bicycle sharing program, the largest in the nation. As in many other cities, helmet use was made optional, in part to encourage greater participation. But a look at the statistics suggests that riding without a helmet is not a decision to [][...]
D.2 medicalxpress.com
@@
medicalxpress.com and healthed.com.au republished Stanford University Medical Center People suffering from a debilitating and often discounted disease known as chronic fatigue syndrome may soon have something they’ve been seeking for decades: scientific proof of their ailment. Researchers at the Stanford University School of Medicine have created a blood test that can flag the disease, which currently lacks a standard, reliable diagnostic test. “Too often, this disease is categorized as imaginary,” said Ron Davis, PhD, professor of biochemistry and of genetics. When individuals with chronic fatigue syndrome seek help from a doctor, they may undergo a series of tests that check liver, kidney and heart function, as well as blood and immune cell counts, Davis said. [][...]
@@
medicalxpress.com and hamodia.com (unavailable) republished The Associated Press Since late last year, people in the central Chinese city of Wuhan have been infected with a viral pneumonia whose cause was unknown. The outbreak raised the specter of another SARS epidemic, which killed hundreds in 2002 and 2003. A preliminary investigation has now identified the respiratory disease as a new type of coronavirus, Chinese state media reported Thursday, citing scientists handling the investigation. As of Sunday, local authorities reported 59 people with the illness. Seven were in critical condition, while the rest were stable. Eight were discharged Wednesday night after they didn’t exhibit any more symptoms for several days. [][...]
D.3 stackexchange.com
@@
Quote found on www-cs-faculty.stanford.edu matches quoted answer [][...] Many readers are no doubt thinking, ”Why does Knuth replace MIX by another machine instead of just sticking to a high-level programming language? Hardly anybody uses assemblers these days.” Such people are entitled to their opinions, and they need not bother reading the machine-language parts of my books. But the reasons for machine language that I gave in the preface to Volume 1, written in the early 1960s, remain valid today: [][...]
@@
Poetry found on en.wikiquote.org matches quoted answer [][...] And the Raven, never flitting,
@@
That lies floating on the floor, Shall be lifted—nevermore. [][...]
Appendix E Additional experimental results
E.1 Top 50 filtered URLs
@@ E.2 More results
Appendix F NYTimes Memorization Experiments
@@
F.1 Benchmark Creation Prompts
@@
F.2 Additional Results
We present memorization quantification results in Table 7, Table 8, and Table 9 for 256, 128, and 512 prompt lengths, respectively. We further added BLEU and 4GP metrics, where 4GP measures word-level 4-gram overlap precision. For each prompt length, we filter those articles with less than the prompt length since there is no completion for this kind of prefix. The evaluation results on the 1T token pretrained model from section 5.3 can be seen in Table 7. For all the generations, we used temperature 0.0 (greedy decoding).
To verify the validity of the NYT-MCQ, we performed a sanity check using DeepSeek-R1. When the model was supplied with the relevant article at inference, it attained a 99.0% accuracy. When the article was withheld, the model achieved 91.7% accuracy. For additional comparison, we tested GPT-4o on NYT-MCQ and observed 89.8% accuracy.
LCCS $\downarrow$ BLEU $\downarrow$ NYT-MCC $\downarrow$ NYT-MCQ $\uparrow$
Qwen2.5-0.5B 19.51 0.71 42.8 24.4
Qwen2.5-1.5B 24.35 1.18 52.1 30.7
Qwen2.5-3B 27.13 1.54 56.4 31.3
Qwen2.5-7B 30.99 1.95 63.8 36.4
Llama-3.2-1.2B 20.71 0.58 49.9 24.8
Llama-3.2-3B 22.75 0.70 57.6 30.1
PT + CD (Non-compl., 100B) 23.11 0.63 50.6 28.9
PT + CD (Compl., 100B ) 21.27 0.51 48.6 27.5
PT + CD (Compl., Decontaminated with News, 100B) 20.28 0.53 48.0 28.6
PT + CD (Compl. with News Removed, 100B) 21.06 0.55 48.9 27.3
PT (Compl., 90B) + CD (Non-compl., 10B) 21.74 0.55 50.4 26.9